OpenStudy (dls):

Integrate :

OpenStudy (dls):

\[\LARGE \int\limits_{0}^{\frac{\pi}{3}} \frac{\sec x \tan x }{1+\sec^2 x}dx\]

OpenStudy (dls):

Attempt: Let secx=t \[\LARGE \int\limits\limits_{0}^{\frac{\pi}{3}} \frac{\cancel{\sec x \tan x }}{1+t^2 } \times \frac{dt}{\cancel{\sec x \tan x}}\] \[\LARGE => \tan^{-1} (t)=\tan^{-1}(\sec x)\]

OpenStudy (psymon):

Thats right.

OpenStudy (dls):

Final answer is not :(

OpenStudy (dls):

\(\LARGE \tan^{-1}(2)\) is what I'm getting.

OpenStudy (psymon):

Well your initial answer without the limits applied was correct O.o

OpenStudy (dls):

but since domain of tan^{-1}x is -pi/2 to pi/2 so..we need some changes here

OpenStudy (dls):

wait :|

OpenStudy (dls):

\[\LARGE \tan^{-1}(\sec{\frac{\pi}{3}})-\tan^{-1}(\sec0)=>\tan^{-1}(2)-\frac{\pi}{4}\]

OpenStudy (psymon):

Yeah, thats what I get.

OpenStudy (dls):

\[\LARGE \int\limits_{0}^{\frac{\pi}{2}} \frac{\cos x}{(1+\sin x)(2+\sin x)}dx\] another problem i am stuck on :(

OpenStudy (psymon):

Did ya try partial fractions?

OpenStudy (dls):

didn't work

OpenStudy (dls):

Tried in a different way. Let sinx=t \[\LARGE \int\limits\limits_{0}^{1} \frac{dx}{(1+t)(2+t)}dt\]

OpenStudy (dls):

now we can use partial

OpenStudy (psymon):

Yep, that gets the correct answer it looks like.

OpenStudy (dls):

A=1 and B= -1 ?

OpenStudy (psymon):

Yep

OpenStudy (dls):

\[\LARGE \int\limits_{0}^{1} \frac{1}{1+t}-\frac{1}{2+t}dt\]

OpenStudy (psymon):

Right.

OpenStudy (dls):

\[\LARGE \log|1+t|-\log|2+t| =>\log (\frac{|1+t | } { |2+t|})\]

OpenStudy (psymon):

yep yep, now back substitute and apply limits.

OpenStudy (dls):

\[\LARGE \log(\frac{2}{3})-\log(\frac{1}{2})=>\log(\frac{4}{3})\]

OpenStudy (psymon):

tada! : )

OpenStudy (psymon):

Yeah, youre awesome with these, once you get the right idea you're golden : )

OpenStudy (dls):

yep :D

OpenStudy (psymon):

good job xD

OpenStudy (dls):

thanks for motivating o.O

OpenStudy (psymon):

Hey, Ive seen ya work, you do a good job, absolutely :3

OpenStudy (dls):

ty :)

OpenStudy (psymon):

:3